9x2 + 12x + 4 = 0
[ Tengo que resolver esta ecuacion cuadratica e indicar las caracteristicas de la ecuacion y cual metodo se utilizo para resolverla. ]
Me podrian ayudar por favor!?
se los agradeceria un monton!!

Answer by TimothyLamb(4379)

9x^2 + 12x + 4 = 0
---
the above quadratic equation is in standard form, with a=9, b=12, and c=4
---
to solve the quadratic equation, plug this:
9 12 4
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has only one root:
x = -0.666666667 = -2/3
---
the quadratic vertex is a minimum at: ( x= -2/3, y= 0 )
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php